OpenStudy (anonymous):

Cyclist rode first 36 mile portion of his race at constant speed. For 27 cooldown portion he reduced his speed by 3 miles per hour. Each portion took same time, find cyclist speed during first portion and speed during cooldown. Show me how to set it up? Thanks.

OpenStudy (king):

let constant speed be v then, t=36/v fr 1st portion

OpenStudy (anonymous):

36/x-3=27/x-3 Is this right?

OpenStudy (king):

fr 2nd portion let speed be v/3 so, t=27/(v-3)

OpenStudy (king):

no its 36/x=27/x-3

OpenStudy (anonymous):